OpenStudy (anonymous):

the integral of 6x^2(e^((-x)^3)) from one to infinity what does it converge at? and how do you get the answer?

OpenStudy (goten77):

no dx????

zepdrix (zepdrix):

\[\Large\bf \int\limits_{x=1}^{\infty} \;e^{\color{#DD4747}{(-x)^3}}\color{#008353 }{6x^2\;dx}\]Making the following u-sub:\[\Large\bf \color{#DD4747}{u=(-x)^3}\]Taking the derivative of our sub and moving some stuff around...\[\Large \bf \color{#008353 }{-2du=6x^2\;dx}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thats all i should need

zepdrix (zepdrix):

k :)

zepdrix (zepdrix):

don't forget to either: ~change the limits of integration so they correspond to u ~or remember to undo your sub after you integrate before plugging in bounds.

OpenStudy (anonymous):

whats the limit of e^-infinity

zepdrix (zepdrix):

Remember your rules of exponents! :) \[\Large e^{-99999}\quad=\quad \frac{1}{e^{99999}}\quad\approx\quad \frac{1}{1\;bajillion}\quad\approx\quad 0\]Does that make sense?

OpenStudy (anonymous):

yes so now i have -2e^(-x^3) from one to infinity so that should be -2e^-1 right

OpenStudy (anonymous):

or -2/e

zepdrix (zepdrix):

\[\Large\bf \lim_{b\to\infty}-2e^{-b^3}-(-2e^{-1})\]I think you missed a minus sign on your lower limit.

OpenStudy (anonymous):

isnt it upper minus lower?

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

thank you soo much

zepdrix (zepdrix):

yay gj \c:/ np!